OpenStudy (anonymous):

How do you solve this linear equation. 5x+7=12?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Can sombody help me solve this problem?

OpenStudy (anonymous):

Anybody? How do you solve this linear equation. 5x+7=12?

OpenStudy (campbell_st):

start by subtracting 7 from both sides of the equation.... what do you get.? 5x + 7 = 12 -7 -7 -------------- ?

OpenStudy (anonymous):

5 @campbell_st

OpenStudy (campbell_st):

well nearly you get 5x = 5 next step is to divide both sides of the equation by 5 \[\frac{5x}{5} = \frac{5}{5}\] so what's the value of a single x...?

OpenStudy (anonymous):

x=1 @campbell_st

OpenStudy (campbell_st):

yes, thats it...

OpenStudy (anonymous):

Thanks @campbell_st